AAVSOnet Status, 2022-09-14

Wed, 09/14/2022 - 12:19

The pipeline/Amazon Cloud transition has stabilized and has been running smoothly, with Cliff doing a great job in making the transition seem seamless.  He is currently working on removing most of the ancillary Fortran code that is used for generating the Epoch Photometry Database, replacing it with a more modern Sextractor/Python version.  George has been working on improving the real-time web-based engineering data.  Ken has been working with the researchers to add targets to the queue, adjust exposure times, etc., and has also been representing AAVSOnet in the monthly section leader meetings.

BSM_Berry.  Operational, just constrained by weather.

BSM_Hamren.  Still offline.  Gary Walker and Josh Walawender are working on bringing this system back to life.

BSM_NH2.  Operational.

BSM_NM.  Operational.

BSM_S.  Offline, waiting for completion of the enclosure.

BSM_TX. Operational.

MPO61. Operational, with weather constraints.

OC61.  Operational.

SRO50.  Offline.  Its new home in Red Hill, NM is nearing completion, and the telescope will be taken there in November by Walt Cooney.

TMO61.  Offline, with computer issues.  NMSU is upgrading the control computer to one running Win11 and some drivers aren't working yet.